OpenStudy (nicia99):

Which is not a power of 10? ten raised to the zero point five power 0.001 100,000

hartnn (hartnn):

" ten raised to the zero point five power " has the power as 0.5 not 10

OpenStudy (agent47):

They are all powers of 10 though.

OpenStudy (nicia99):

Thanks :)

OpenStudy (agent47):

10^x where x is some number will be a power of 10.

OpenStudy (agent47):

the first is 10^(0.5), next is 1/1000 = 1/10^3 = 10^(-3)

hartnn (hartnn):

not actually, 10^0.5 is not, its actually sqrt 10

OpenStudy (agent47):

Yes, I know that, but wouldn't that still be considered a 0.5th power of 10?

hartnn (hartnn):

given the choices, thats the best answer as the question assumes the integral power

OpenStudy (agent47):

I kinda disagree, but have no problems with your answer if she got it right. Question is silly though, and I would go for all of the above as the answer.